### v1

# arXiv:2603.01273v1[cond-mat.quant-gas] 01 Mar 2026 thanks: ORCID: 0000-0003-2624-0902

Abstract
We present a geometric framework for 1D quantum fluids across interaction regimes in the Thomas-Fermi limit. Based on the Linearization Principle via the $q$-logarithm, macroscopic density profiles form a discrete hierarchy: the ideal Bose gas ($q=1$), the mean-field Gross-Pitaevskii condensate (q=1q=-1), and the strongly correlated Tonks-Girardeau gas (q=3q=-3). We further derive a universal sound velocity scaling, $c\propto\rho^{(1-q)/4}$. This establishes a non-perturbative link between static geometry and dynamical excitations in many-body systems.
Gross-Pitaevskii equation, Tonks-Girardeau gas, Thomas-Fermi limit, Linearization Principle, Information geometry
pacs:
03.75.Hh, 05.30.Jp, 05.90.+m
I Introduction
The Gross-Pitaevskii equation (GPE) is the standard mean-field model for Bose-Einstein condensates (BECs) [1]. As a nonlinear Schrödinger equation, it successfully captures the macroscopic behavior of quantum fluids within the mean-field approximation, where the inter-particle interaction is effectively described by a cubic nonlinearity. However, the validity of the standard GPE is limited to the weak-coupling regime. In one-dimensional (1D) systems, as the effective interaction strength increases, the gas enters a strongly correlated regime where the mean-field approximation breaks down. The extreme limit of infinite repulsion is the Tonks-Girardeau (TG) gas, where bosons mimic the exclusion statistics of fermions due to impenetrability [2, 3]. The density profile of the TG gas in a harmonic trap follows the Wigner semicircle law, which is fundamentally different from the inverted parabola profile (Thomas-Fermi approximation) of the standard BEC.
The exact microscopic description of 1D interacting bosons with contact interactions is provided by the Lieb-Liniger model [4]. The interaction regime is characterized by the dimensionless parameter $\gamma=m_{0}g/\hbar^{2}\rho$, where $m_{0}$ is the particle mass, $g$ is the interaction strength, and $\rho$ is the 1D density. While the Lieb-Liniger model is exactly solvable via the Bethe ansatz for a homogeneous system, extracting the macroscopic density profile $n(x)$ in the presence of an external harmonic trap typically relies on the Local Density Approximation (LDA) combined with numerical solutions. Conventionally, the weak and strong correlation limits—the mean-field BEC ($\gamma\ll 1$) and the TG gas ($\gamma\gg 1$)—are described by distinct mathematical frameworks: the nonlinear GPE for the former, and the Bose-Fermi mapping for the latter. This raises the question of whether macroscopic density profiles across these disparate correlation regimes can be analytically captured within a single geometric framework, bypassing the need for numerical LDA.
In this Letter, we introduce a geometric framework, the Linearization Principle, to describe these systems in a unified manner. Within this framework, macroscopic density deformations arise from the geometric curvature of the state space. By introducing the $q$-logarithm [5, 6] as a natural coordinate that linearizes the stationary state, we derive a family of analytical density profiles in the Thomas-Fermi limit. This geometric framework reveals a discrete hierarchy governing quantum fluids. We show that a single geometric index $q$ parameterizes the physical regime: $q=1$ yields the ideal Bose gas, q=1q=-1 recovers the Thomas-Fermi profile of the GPE, and q=3q=-3 yields the fermionized Wigner semicircle law of the TG gas. Furthermore, we demonstrate that this geometric structure extends to dynamical excitations, yielding the sound velocity scaling $c\propto\rho^{(1-q)/4}$ across the interaction regimes.
II Theory
II.1 The Nonlinear Schrödinger Equation and the Thomas-Fermi Limit
The starting point of our discussion is the standard Gross-Pitaevskii equation (GPE), which describes the ground state of a Bose-Einstein condensate in the mean-field approximation:
$i\hbar\frac{\partial\psi}{\partial t}=-\frac{\hbar^{2}}{2m_{0}}\nabla^{2}\psi+V_{\text{ext}}(x)\psi+g|\psi|^{2}\psi,$ (1)
where $g$ represents the interaction strength. For a large number of particles with repulsive interactions, the interaction energy dominates over the kinetic energy (quantum pressure) associated with the spatial variations of the density. Neglecting the kinetic energy term $\frac{\hbar^{2}}{2m_{0}}\nabla^{2}\psi$ leads to the Thomas-Fermi (TF) approximation. Our goal is to derive the macroscopic density profile of this limit—and its generalizations for strong coupling regimes—from a geometric principle.
II.2 The Linearization Principle
We introduce the Linearization Principle. The core of this principle is the mathematical property that a nonlinear differential equation of the form $dy/dx=y^{q}$ can be transformed into a linear equation:
$\frac{dy}{dx}=y^{q}\quad\Longrightarrow\quad\frac{d\ln_{q}y}{dx}=1,$ (2)
where the $q$-logarithm is defined as $\ln_{q}x:=(x^{1-q}-1)/(1-q)$. This $q$-logarithm serves as a natural coordinate that maps the nonlinear stationary state into a linear form.
In the TF limit, where the spatial second derivative is neglected, the quantum fluid reaches a local mechanical equilibrium. Physically, this means that the internal generalized force (the gradient of the local chemical potential driven by interactions) must balance the external trapping force, $\nabla V_{\text{ext}}(x)$. Within our geometric framework, this fundamental force balance is captured by a linear response relation in the $q$-logarithmic scale:
$\frac{d}{dx}\ln_{q}\psi(x)=-\beta\frac{dV_{\text{ext}}(x)}{dx},$ (3)
where $\beta$ is a positive constant related to the interaction strength and the chemical potential.
The physical origin of this linear response in the $q$-logarithmic scale is rooted in the information geometry of strongly correlated systems [7, 8]. In standard statistical mechanics, the ideal gas ($q=1$) is governed by the additive Boltzmann-Gibbs entropy, leading to a linear differential equation in the standard logarithmic scale. However, the presence of mean-field interactions or hard-core repulsions deforms the accessible state space. The index $q$ quantifies this geometric deformation. By transforming the system into the $q$-logarithmic coordinate, the complex nonlinear correlations are unfolded into a mathematically flat (linear) space, allowing us to describe the macroscopic density profile without relying on perturbative expansions.
II.3 Geometric Wavefunction and Density Profile
Integrating Eq. (3) yields the $q$-Gaussian effective wavefunction:
$\psi(x)\propto\left[1-(1-q)\beta V_{\text{ext}}(x)\right]_{+}^{\frac{1}{1-q}},$ (4)
where $X]_{+}=\max(X,0)$ denotes the cut-off, representing the physical boundary of the quantum gas cloud. The macroscopic particle density distribution, $n(x)=|\psi(x)|^{2}$, is given by:
$n(x)=n(0)\left[1-(1-q)\beta V_{\text{ext}}(x)\right]_{+}^{\frac{2}{1-q}}.$ (5)
II.4 Derivation of the Standard GPE (q=1q=-1)
[…]
$n(x)=n(0)\left[1-2\beta V_{\text{ext}}(x)\right]_{+}.$ (6)
This is the inverted parabola profile known as the Thomas-Fermi approximation, which describes the ground state of a standard BEC.
[…]
$\beta V_{\text{ext}}(x)\psi+\frac{1}{2}\psi^{3}=\mu^{\prime}\psi,$ (7)
where $\mu^{\prime}$ is an integration constant. Comparing this to the stationary GPE in the TF approximation ($V_{\text{ext}}\psi+g|\psi|^{2}\psi=\mu\psi$), we identify the mapping between the geometric parameter $\beta$ and the physical parameters. The chemical potential $\mu$ is uniquely determined by the normalization condition $\int|\psi(x)|^{2}dx=N$ and encapsulates the microscopic interaction strength $g$. This confirms that the Thomas-Fermi regime of the standard GPE is a mathematical consequence of the q=1q=-1 geometry.
III Geometric Hierarchy to the Tonks-Girardeau Limit
Having identified q=1q=-1 as the standard mean-field regime, we now explore the strong-coupling limit.
III.1 Derivation of the Semicircle Law (q=3q=-3)
In the limit of infinite repulsion, a 1D Bose gas becomes fermionized, forming a Tonks-Girardeau (TG) gas [2, 3]. For a system confined in a harmonic trap, $V_{\text{ext}}(x)\propto x^{2}$, the ground-state density profile is described by the Wigner semicircle law:
$n_{\text{TG}}(x)=n(0)\left[1-\left(\frac{x}{R}\right)^{2}\right]_{+}^{\frac{1}{2}},$ (8)
where $R$ is the cloud radius.
[…]
III.2 Discrete Physical Hierarchy
These results reveal a discrete hierarchy of integer $q$-values corresponding to fundamental macroscopic states (see Fig. 1):
  • $q=1$: Ideal Bose gas (Gaussian profile).
  • q=1q=-1: Standard BEC in the TF limit (Inverted parabola profile).
  • q=3q=-3: Tonks-Girardeau gas (Wigner semicircle profile).
Figure 1: Normalized density profiles $n(x)/n(0)$ in a harmonic trap derived from the Linearization Principle for different values of the geometric index $q$. The profiles demonstrate the discrete hierarchy from the ideal Bose gas ($q=1$, dashed line), to the mean-field BEC in the Thomas-Fermi limit (q=1q=-1, solid line), and the strongly correlated Tonks-Girardeau gas (q=3q=-3, dash-dotted line).
IV Theoretical Consistency: Connection to Nonlinear Diffusion and Geometry
In this section, we establish the thermodynamic and dynamical consistency of the geometric hierarchy by mapping our index $q$ to the polytropic index $m$ of the nonlinear diffusion equation.
IV.1 Correspondence with the Porous Medium Equation
In the hydrodynamic limit, the macroscopic density dynamics can be modeled by the Porous Medium Equation (PME) [9]:
$\frac{\partial\rho}{\partial t}=\nabla\cdot\left[D\nabla(\rho^{m})+\rho\nabla V_{\text{ext}}\right],$ (9)
whose stationary solution yields the density profile:
$\rho(x)=\rho(0)\left[1-\frac{m-1}{mD}V_{\text{ext}}(x)\right]_{+}^{\frac{1}{m-1}}.$ (10)
Comparing this spatial exponent with our geometric density exponent 2/(1q)2/(1-q), we immediately obtain the fundamental mapping m=(3q)/2m=(3-q)/2.
IV.2 Equations of State and Thermodynamics
In the hydrodynamic framework, the PME index $m$ is related to the polytropic equation of state, $P\propto\rho^{m}$. Applying our geometric mapping m=(3q)/2m=(3-q)/2 yields:
  • Standard BEC (q=1q=-1): The mapping yields $m=2$, corresponding to the equation of state $P\propto\rho^{2}$. This is the pressure-density relation for a mean-field BEC arising from two-body contact interactions.
  • Tonks-Girardeau Gas (q=3q=-3): The mapping yields $m=3$, corresponding to the equation of state $P\propto\rho^{3}$. This recovers the degeneracy pressure exerted by a 1D gas of hard-core bosons (fermionization).
Thus, the geometric index $q$ encodes the thermodynamics of the quantum fluid across disparate interaction regimes.
IV.3 Geometric Scaling of the Sound Velocity
The geometric framework also governs the collective dynamical excitations. In the hydrodynamic regime, the local speed of sound $c$ is determined by the thermodynamic relation $c=\sqrt{\partial P/\partial(m_{0}\rho)}$, where $m_{0}$ is the particle mass. Using the geometric equation of state $P\propto\rho^{(3-q)/2}$, the sound velocity universally scales with the density as:
$c\propto\sqrt{\frac{\partial}{\partial\rho}\left(\rho^{\frac{3-q}{2}}\right)}\propto\rho^{\frac{1-q}{4}}.$ (11)
This single geometric scaling law captures the fundamental dynamic responses of the system:
  • For the mean-field BEC (q=1q=-1), Eq. (11) yields $c\propto\rho^{1/2}$, recovering the Bogoliubov sound velocity.
  • For the Tonks-Girardeau gas (q=3q=-3), it yields $c\propto\rho^{1}$, matching the Fermi velocity of a 1D ideal Fermi gas.
Therefore, the geometric parameter $q$ not only determines the static macroscopic density profile but also dictates the dynamical response and collective excitations of the quantum many-body system.
V Experimental Signatures and the Crossover Regime
The analytical geometric hierarchy derived here is amenable to experimental verification using ultracold atomic gases. By tuning the effective 1D interaction strength $\gamma$ via confinement-induced resonances, experimentalists can sweep the system across the correlation regimes. Our framework establishes q=1q=-1 and q=3q=-3 as the mathematical limits (geometric fixed points) of this tunable system. In these extreme limits, the in situ density profiles can be observed to confirm the $q$-Gaussian forms, and the corresponding sound velocity scaling $c\propto\rho^{(1-q)/4}$ can be independently probed.
In the intermediate crossover regime ($\gamma\sim 1$), the macroscopic density profile smoothly interpolates between the inverted parabola and the Wigner semicircle, as described by the Lieb-Liniger equation of state combined with the Local Density Approximation (LDA). Within our geometric framework, this crossover regime corresponds to a transition phase where the macroscopic state space is not globally characterized by a single integer index. Therefore, rather than being mere asymptotic approximations, the precise integer indices q=1q=-1 and q=3q=-3 physically emerge as the fundamental geometric fixed points that bracket the many-body physics of 1D quantum fluids.
VI Conclusion
In this paper, we have proposed a geometric approach to the density profiles of 1D quantum fluids based on the Linearization Principle. By introducing the $q$-logarithm, we derived macroscopic stationary density profiles in the Thomas-Fermi limit without relying on perturbation theory. We identified a discrete geometric hierarchy: the index q=1q=-1 corresponds to the standard mean-field BEC (parabolic profile), whereas the index q=3q=-3 yields the density profile of the Tonks-Girardeau gas (Wigner semicircle law in a harmonic trap). We established the thermodynamic consistency of this hierarchy by mapping the geometric index $q$ to the polytropic index $m$ via the relation m=(3q)/2m=(3-q)/2, confirming the $P\propto\rho^{2}$ and $P\propto\rho^{3}$ equations of state. Furthermore, this static geometric structure governs the dynamical excitations, yielding the sound velocity scaling $c\propto\rho^{(1-q)/4}$ across the interaction regimes.
The analytical geometric hierarchy is amenable to experimental verification. By tuning the effective 1D coupling strength, the system can be swept across the interaction regimes. Observing the in situ density profiles $n(x)$ and the density-dependent sound velocity provides an experimental test of the fundamental geometric boundaries, q=1q=-1 and q=3q=-3.
In conclusion, we have shown a geometric progression underlying the density distributions of 1D quantum fluids. The distinct regimes of the ideal Bose gas, the mean-field condensate, and the fermionized TG gas emerge as geometric fixed points, governed by q=1,1q=1,-1, and 3-3, respectively. Our Linearization Principle provides an analytical framework for analyzing macroscopic states of strongly correlated quantum systems, connecting quantum mechanics, nonlinear statistical physics, and information geometry [8].